Calculate the magnitude of the magnetic field from a circular loop of wire of radius 0.20 m, carrying a current of 2.4 A, and with 300 turns of wire at a distance of 2.0 m away from the loop along the axis of the loop.
Answer #1

B = N*mue*I*R^2/(2*(R^2+x^2)^3/2

= 300*4*pi*10^-7*2.4*0.2^2/(2*(0.2^2+2^2)^1.5

= 2.23*10^-6 T